# PanoNormal: Monocular Indoor 360° Surface Normal Estimation

### v1

@@ 5 Conclusion
We present, PanoNormal, a deep architecture designed for single 360° surface normal estimation. Our approach leverages the combined strengths of convolutional neural Networks for the effective capture of local details and spherical vision transformers for global dependency with distortion awareness. Our experiments demonstrate the effectiveness of our framework. PanoNormal exhibits significant advances over state-of-the-art methods across five popular 360° panoramic datasets. Additionally, PanoNormal demonstrates enhanced generalizability compared to previous works, and our ablation studies reinforce the efficacy of our proposed architecture and indicate its potential applicability in real-world scenarios. In the future, we will investigate the utility of our deep architecture that combines spherical transformers and convolutional operations in other 360° regression tasks.

### v2

@@ 7 Conclusion
We present PanoNormal, a deep architecture designed for single 360° surface normal estimation. Our approach leverages the combined strengths of convolutional neural networks for effectively capturing local details and spherical vision transformers for global dependency with distortion awareness. Experiments demonstrate PanoNormal’s significant advances over state-of-the-art methods across five popular 360° panoramic datasets. Additionally, PanoNormal demonstrates enhanced generalizability compared to previous works, with our ablation studies further validating the effectiveness of our proposed architecture. Our findings also reveal that models originally designed for 360° depth estimation are not directly transferable to surface normal prediction, underscoring the necessity of tailored architectural design. This makes it a reliable tool for tasks that demand precise environmental understanding.
Acknowledgements
This research was supported by the Marsden Fund Council managed by the Royal Society of New Zealand (No. MFP-20-VUW-180) and Tiandi Science and Technology Co. through the Technology Collaboration Project (No. 2025-TD-CXY003).
[…]
The authors declare that they have no conflict of interest.
[…]
Kun Huang: Conceptualization, Paper writing, Method design, Experiment design & conduct. Jianwei Yang: Experiment design, Data processing. Tielin Zhao: Experiment design, Data processing. Lei Ji: Experiment design, Data processing. Songyang Zhang: Experiment design, Data processing. Fang-Lue Zhang: Method design, Experiment design and Paper writing. Neil Dodgson: Method design, Experiment design and Paper writing.